# Focus On This, Not That! Steering LLMs With Adaptive Feature Specification

### v1

@@ A.6 Spurious Sentiment (SS) Dataset
The sentiment feature still remains causal within the augmented SS dataset, fully predicting the label $Y$ for each dataset example.
@@
A.7 Spurious NLI dataset (SMNLI)
@@
  • $C$ NLI relationship between a premise and hypothesis pair, the causal feature within this task, sampled from the original dataset.
@@
$C$$\sim\begin{cases}\hat{C}(S)&\text{ if }U_{C}=1;\\ R\sim U(\mathcal{C}\setminus\hat{C}(S))&\text{ if }U_{C}=0;\end{cases}\text{ where }\mathcal{C}=\{0,1,2\}\text{ is the entailment label set};$ (48)
@@
$Y$ $=S.$ (50)
Here, $\texttt{Cat}(\mathcal{S})$ is a uniform categorical distribution over spurious feature values (here the underlying genre of the premise-hypothesis pairs). Furthermore, we define $\hat{C}(S)$ to be the entailment label that a particular value of $S$ is spuriously correlated with by design. Moreover, $p_{\mathcal{D}}(\cdot|C,S)$ is the conditional distribution over the dataset examples (premise-hypothesis pairs) that have NLI relationship $C$ and genre $S$.
We restrict the genres within the model to $S\in\{\textit{slate},\textit{government},\textit{fiction},\textit{travel}\}$, a subset of the genres of the training set. When creating a distribution shifted test set, we restrict the genres to $S\in\{\textit{facetoface},\textit{nineeleven},\textit{verbatim}\}$. The specific spurious correlations between genres and labels $Y$ are chosen to be:
  1. 1.
    $\hat{C}(\textit{slate})=0$;
  2. 2.
    $\hat{C}(\textit{government})=2$;
  3. 3.
    $\hat{C}(\textit{fiction})=1$;
  4. 4.
    $\hat{C}(\textit{travel})=0$;
  5. 5.
    $\hat{C}(\textit{facetoface})=2$;
  6. 6.
    $\hat{C}(\textit{nineeleven})=0$;
  7. 7.
    $\hat{C}(\textit{verbatim})=1$.
@@
In this way we generate spurious correlations within the dataset through sub-sampling to induce spurious correlations between $S$ and $Y$. In this case, for a dataset example $X$, the spurious correlation between $S=s$ and $\hat{C}(s)$ is given as
$\mathbb{P}(Y=\hat{C}(s)|X,S=s,s\in X)=\rho_{\text{spurious}}.$ (51)
This follows immediately form the sub-sampling process described above. This then once again aligns with the definition of $\rho_{\text{spurious}}$ in Equation 4 and justifies the choice of notion in the above generative process.
@@
A.8 Spurious HANS dataset (SHANS)
@@
  • $C$ NLI relationship between a premise and hypothesis pair, the causal feature within this task, sampled from the original dataset.
@@
$C$$\sim\begin{cases}\hat{C}(S)&\text{ if }U_{C}=1;\\ R\sim U(\mathcal{C}\setminus\hat{C}(S))&\text{ if }U_{C}=0;\end{cases}\text{ where }\mathcal{C}=\{0,1\}\text{ is the entailment label set};$ (58)
@@
$Y$ $=S.$ (60)
Here, $\texttt{cat}(\mathcal{S})$ is a uniform categorical distribution over $\mathcal{S}$ which effectively selects the precence of exactly one of the three spurious feature heuristics. We define $\hat{C}(S)$ to be the entailment label that a particular value of $S$ is spuriously correlated with by design. Moreover, $p_{\mathcal{D}}(\cdot|C,S)$ is the conditional distribution over the dataset examples (premise-hypothesis pairs) that have NLI relationship $C$ and the presence of spurious heuristics $S$.
We consider the presence of each feature to be separate binary spurious features. The specific spurious correlations between heuristics and labels $Y$ are chosen to be:
  1. 1.
    $\hat{C}(S_{\text{lex.}})=0$;
  2. 2.
    $\hat{C}(S_{\text{sub.}})=0$;
  3. 3.
    $\hat{C}(S_{\text{const.}})=1$;
In this way we generate spurious correlations within the dataset through sub-sampling to induce spurious correlations between the heuristics and $Y$. In this case, for a dataset example $X$, the spurious correlation between the presence of a particular heuristic $S=s$ and $\hat{C}(s)$ is given as
$\mathbb{P}(Y=\hat{C}(s)|X,S=s,s\in X)=\rho_{\text{spurious}}.$ (61)
This follows immediately form the sub-sampling process described above. Once again, this aligns with the definition in Equation 4 which again justifies the introduced notation.
@@ A.9 FIT on SHANS
Figure 14: SHANS focus accuracies ($\uparrow$). Focus accuracy ($\mathcal{A}_{\text{focus}}$) of Llama-3.1-8B-Instructafter FIT on the SHANS dataset. Here, $C$ refers to the causal feature (logical relationship between premise and hypothesis) and $S$ the spurious feature (heuristic used)

### v2

@@ A.9 Spurious Sentiment (SS) Dataset
The sentiment feature still remains core within the augmented SS dataset, fully predicting the label $Y$ for each dataset example.
@@
A.10 Spurious NLI dataset (SMNLI)
@@
  • $C$ NLI relationship between a premise and hypothesis pair, the core feature within this task, sampled from the original dataset.
@@
$C$ =$=\begin{cases}y_{S}&\text{ if }U_{C}=1;\\ \sim\texttt{Unf}(\text{Val}(Y)\setminus\{y_{S}\})&\text{ if }U_{C}=0;\end{cases};$ (39)
@@
$Y$ $=C.$ (41)
Here, $\texttt{Cat}(\mathcal{S})$ is a uniform categorical distribution over spurious feature values (here the underlying genre of the premise-hypothesis pairs). Furthermore, we define $y_{S}$ to be the NLI label that a particular value of $S$ is spuriously correlated with by design. Moreover, $p_{\mathcal{D}}(x|C,S)$ is the conditional distribution over the dataset examples (premise-hypothesis pairs) that have NLI relationship $C$ and genre $S$.
We restrict the genres within the model to $S\in\{\textit{slate},\textit{government},\textit{fiction},\textit{travel}\}$, a subset of the genres of the training set. When creating a distribution shifted test set, we restrict the genres to $S\in\{\textit{facetoface},\textit{nineeleven},\textit{verbatim}\}$. The specific spurious correlations between a genre $s$ and a label $y_{s}$ are chosen to be: $y_{\textit{slate}}=0$; $y_{\textit{government}}=2$; $y_{\textit{fiction}}=1$; $y_{\textit{travel}}=0$; $y_{\textit{facetoface}}=2$; $y_{\textit{nineeleven}}=0$; $y_{\textit{verbatim}}=1$.
@@
In this way we generate spurious correlations within the dataset through sub-sampling to induce spurious correlations between $S$ and $Y$.
We show that the notion of $\rho_{\text{spurious}}$ in Equation 38 aligns with the notation in Definition 1.
Proposition 4.
From the SCEs described above, we have that
$\mathbb{P}(Y=y_{s}\mid S=s)=\rho_{\text{spurious}}\;.$ (42)
for all spurious feature values $s$.
Proof.
This is clear considering Equation 39, where $\rho_{\text{spurious}}$ influences the chance that we sample $y_{s}$, i.e. the label spuriously correlated with feature value $s$. ∎
@@
Independence conditions during training for FIT. As specified in Section A.1, we would like to have that $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ during training so that models trained via FIT can effectively learn to leverage focus instructions to make predictions based on specified features, where, again, $Y_{S}$ is the label spuriously associated to spurious feature value $S$. The results below give sufficient conditions for this to occur with respect to the DGP described in Figure 14.
Proposition 5.
Assuming the SCEs given above and the corresponding DGP described in Figure 14, if we have a uniform label distribution $p(Y=y)$, and have that $\rho_{\text{spurious}}=1/3$ in the SMNLI training set, then we have that $Y\perp\!\!\!\perp S$.
Proof.
From Proposition 4, we have that $\mathbb{P}(Y=y_{s}\mid S=s)=1/3$ when $\rho_{\text{spurious}}=1/3$ for any spurious feature value $s\in\text{Val}(S)$. Moreover, we have that from Equation 39
$\mathbb{P}(Y=y\mid S=s)$ $=\mathbb{P}(C=y\mid S=s)$ (43)
$=\underbrace{\mathbb{P}(U_{C}=1)}_{=\frac{2}{3}}\underbrace{\mathbb{P}(C=Y|S=s)}_{=\frac{1}{2}}$ (44)
$=\frac{1}{3},$ (45)
for each $y\in\text{Val}(Y)\setminus\{y_{s}\}$.
Using this alongside the assumed balanced label distribution, $\mathbb{P}(Y=y)=1/3$ for $y\in\text{Val}(Y)$, implies that we have $\mathbb{P}(Y=y|S=s)=\mathbb{P}(Y=y)$ for all $y\in\text{Val}(Y)$ and $s\in\text{Val}(S)$, which gives that $Y\perp\!\!\!\perp S$. ∎
Proposition 6.
Assuming the SCEs given above and the corresponding DGP described in Figure 14, if we have a uniform label distribution $p(Y=y)$, and have that $\rho_{\text{spurious}}=1/3$ in the SMNLI training set, then we have that $Y_{S}\perp\!\!\!\perp C$.
Proof.
[…]
First, using that $\rho_{\text{spurious}}=1/3$, we have from Equation 39 that
$\mathbb{P}(C=c\mid S=s)$ =$=\begin{cases}1/3&\text{if }c=f(s),\\ 2/3\cdot 1/2&\text{else.}\end{cases}$ (46)
$=\frac{1}{3},$ (47)
[…]
Now we consider the marginal distribution of C. Let $|\text{Val}(S)|=k$ be the number of spurious feature values. Using the previous result, we have that:
$\mathbb{P}(C=c)$ $=\sum_{s}\underbrace{\mathbb{P}(C=c\mid S=s)}_{=\frac{1}{3}}\underbrace{\mathbb{P}(S=s)}_{\frac{1}{k}}$ (48)
$=\sum_{s}\frac{1}{3k}$ (49)
$=\frac{1}{3}.$ (50)
Therefore, we have that $\mathbb{P}(C=c\mid S=s)=\mathbb{P}(C=c)$, for all $c\in\text{Val}(C)$ and $s\in Val(S)$, which then implies that $S\perp\!\!\!\perp C$. Therefore, using that the spurious label $Y_{S}$ is a deterministic function of $S$, we have that $Y_{S}\perp\!\!\!\perp C$ under the assumptions within the proposition. ∎
A.11 Spurious HANS dataset (SHANS)
@@
  • $C$ NLI relationship between a premise and hypothesis pair, the core feature within this task, sampled from the original dataset.
@@
$C$$\sim\begin{cases}y_{S}&\text{ if }U_{C}=1;\\ \sim U(\text{Val}(Y)\setminus\{y_{S}\})&\text{ if }U_{C}=0;\end{cases};$ (57)
@@
$Y$ $=C.$ (59)
Here, $\texttt{Cat}(\mathcal{S})$ is a uniform categorical distribution over $\mathcal{S}$ which effectively selects the precence of exactly one of the three spurious feature heuristics. We define $y_{S}$ to be the NLI label that a particular value of $S$ is spuriously correlated with by design. Moreover, $p_{\mathcal{D}}(x|C,S)$ is the conditional distribution over the dataset examples (premise-hypothesis pairs) that have NLI relationship $C$ and the presence of spurious heuristics $S$.
We consider the presence of each feature to be separate binary spurious features. The specific spurious correlations between heuristics and labels $Y$ are chosen to be: $y_{S_{\text{lex.}}=1}=0$; $y_{S_{\text{sub.}}=1}=0$; $y_{S_{\text{const.}}=1}=1$.
In this way we generate spurious correlations within the dataset through sub-sampling to induce spurious correlations between the heuristics and $Y$.
Transferred results from SMNLI. As we use the same DGP as for the SMNLI dataset described in Section A.11, all of the results that we have proven for SMNLI, translate to the SHANS dataset. In particular, we have that $\rho_{\text{spurious}}$ aligns with the notation in Definition 1, and that we have $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ under the assumptions of balanced label distributions and $\rho_{\text{spurious}}=1/2$ within the training set used for FIT training.
@@ A.12 FIT on SHANS
Figure 18: SHANS focus accuracies ($\uparrow$). Focus accuracy ($\mathcal{A}_{\text{focus}}$) of Llama-3.1-8B-Instructafter FIT on the SHANS dataset. Here, $C$ refers to the core feature (logical relationship between premise and hypothesis) and $S$ the spurious feature (heuristic used)